1) y/4 - 1 = y/5
2) 2y-3/9 + y-3/2 = y+5/6 - 1
3) A= 1/2(a+b) solve for b

Answer by pushpaharan(47)

1) y/4 - 1 = y/5
y/4-y/5 = 1
5y/20 -4y/20 =1
y/20 =1
y=20
2) 2y-3/9 + y-3/2 = y+5/6 - 1
2(2y-3)/18 + 9(y-3)/18 = 3(y+5)/18 - 18/18
4y - 6 + 9y -15 = 3y+ 15 -18
10y = 18
y =1.8
3) A= 1/2(a+b) solve for b
2A = a+ b
b = 2A-a